### v1

# CARE: a Benchmark Suite for the Classification and Retrieval of Enzymes

Abstract
Enzymes are important proteins that catalyze chemical reactions. In recent years, machine learning methods have emerged to predict enzyme function from sequence; however, there are no standardized benchmarks to evaluate these methods. We introduce CARE, a benchmark and dataset suite for the Classification And Retrieval of Enzymes (CARE). CARE centers on two tasks: (1) classification of a protein sequence by its enzyme commission (EC) number and (2) retrieval of an EC number given a chemical reaction. For each task, we design train-test splits to evaluate different kinds of out-of-distribution generalization that are relevant to real use cases. For the classification task, we provide baselines for state-of-the-art methods. Because the retrieval task has not been previously formalized, we propose a method called Contrastive Reaction-EnzymE Pretraining (CREEP) as one of the first baselines for this task. CARE is available at https://github.com/jsunn-y/CARE/.
1 Introduction
Proteins, which are sequences of amino acid building blocks, are not only integral components of all living organisms, but also important for a myriad of commercial applications spanning from the health domain to the bio-economy. Enzymes are a subclass of proteins that can catalyze chemical reactions, and they have many applications in areas such as bioremediation, plastic degradation, gene editing, and drug synthesis [1, 2, 3].
Identifying the specific chemical reactions that an enzyme is capable of performing (i.e. the enzyme’s function) is a key first step for many applications [4]. For instance, during standard metagenomic analyses, enzyme genes are annotated for their functions, which enables determination of an organism’s critical metabolic pathways and specialization [5, 6, 7]. When applied to chemical synthesis, enzyme annotations are needed to identify catalysts that can replace existing step(s) in drug synthesis procedures–known as retrobiosynthesis [8]. Likewise, enzyme engineering for new-to-nature function [9] involves discovering an enzyme starting point with a desired function before improving activity for that function using protein engineering techniques [10, 11]. Historically, similarity search algorithms, most notably BLAST, have been the most common methods used to assign function to protein sequences [12, 13]. These methods work by finding similar sequences in annotated reference databases, as similar sequences are likely to share function. Still, there is a need for more abundant, high-quality annotations of enzymes and automated workflows to identify enzymes with desired functions.
In recent years, there has been increasing interest in using machine learning (ML) for a broad range of applications related to the functional prediction and design of enzymes [11, 14, 15]. In particular, ML models have emerged to classify protein sequences based on their function, which are reviewed here [16] with some more recent examples listed here [17, 7, 18, 19, 20, 21, 22]. Despite these advancements, there is no standard benchmark or dataset for evaluating computational models for enzyme function prediction. A challenge associated with classification is that a given enzyme is often able to perform multiple reactions [23], and many reactions are not annotated. Moreover, complex tasks such as extrapolation to unannotated reactions have yet to be evaluated.
In this work, we present a benchmark suite for the Classification And Retrieval of Enzymes (CARE). Our contributions can be summarized as: (1) formalizing model evaluation into two tasks which are directly relevant to scientists and engineers: classification of an enzyme by function (Task 1), and retrieval of enzyme sequences based on a reaction (Task 2); (2) curating high-quality and easy-to-use datasets; (3) providing train-test splits that mimic challenging extrapolations in real-world use cases; and (4) benchmarking state-of-the-art models for Task 1, and providing the first baselines for Task 2. Because Task 2 has not been previously explored, we introduce a model called Contrastive Reaction-EnzymE Pretraining (CREEP) to serve as an initial baseline. CREEP can perform contrastive learning across three different modalities (protein, reaction, and textual description), and the learned representations are then used for retrieval. Overall, we anticipate that CARE will be a useful and easy-to-use resource for ML researchers to benchmark their enzyme function prediction models.
2 Related Work
Datasets. Various databases have emerged to help researchers store, share, and identify functionally annotated enzymes. Protein sequence databases such as UniProt [24] and Pfam [25] are catalogues of annotated protein sequences. While most databases reference protein sequences, increasingly, these sequences can be linked to protein structures, either experimentally validated, as in the Protein Data Bank (PDB), or via structural prediction tools [26] in databases such as the AlphaFold Database [27]. BRENDA [28] is a curated database specific for reaction and enzyme sequence information. Rhea [29] consolidates information from BRENDA, and other sources, such as pathway databases including KEGG [30]. There is ongoing work to compile and clean/standardize reactions from multiple databases, namely ECReact [31] and EnzymeMap [32]. Related to these databases, RetroBioCat is a database and retrobiosynthesis planning tool [33, 34], and Selenzyme is tool to retrieve enzymes to perform a target reaction [35].
Protein Benchmarks. Our work takes inspiration from existing protein fitness prediction benchmarks, where fitness is a quantification of some function. TAPE [36] and FLIP [37] evaluate representations from protein language models for the prediction of a broad range of general and specific protein properties (stability, secondary structure, binding etc.) ProteinGym considers sequence variant effect prediction by using likelihoods from these language models [38]. While benchmarks for protein fitness prediction tasks are well defined, there are no standardized benchmarks for protein function prediction, which is more qualitative/categorical rather than quantitative. Here, we focus on proteins that perform chemical reactions (i.e., enzymes).
Classification of Enzyme Function. ML models have emerged to predict the outcomes of enzymatic reactions [39, 40, 41] and for classification of enzyme function. Enzyme function is usually expressed using enzyme commission (EC) numbers, which is a hierarchical scheme for classifying enzyme function into classes (families) and consists of four levels of descriptions (Figure 1A). Some classification models are general protein function prediction models, which encompass all proteins, not just enzymes, such as ProtCNN/ENN [20] and ProteInfer [21]. Many models utilize representations from protein language models [42, 43, 44], and others incorporate protein structure as information, such as DeepFRI [19] and BioCLIP [45]. Methods related to supervised contrastive learning [46] have been particularly useful here, such as CLEAN and HiFi-NN [18, 7], likely by reducing imbalances in the number of sequences representing each EC number. Recently, combining ML with BLAST has enabled retrieval-augmented function prediction in ProtEx [22].
Large Language Models. Recently, there has been an explosion in pretrained models in the biological and chemical domains, particularly large language models (LLMs) [47]. For example, ChatGPT is capable of answering questions related to general scientific knowledge. These language models can be further finetuned for applications such as answering questions about protein sequences (Pika, [48]) and for reaction synthesis planning (ChemCrow, [49]), among others. LLMs present important benchmarks for enzyme functional classification and retrieval, given the widespread adoption of LLMs as science facilitators and their ease of use (in particular webserver-based approaches), compared to domain-specific methods. In this study, we include ChatGPT as a baseline.
Multimodal Contrastive Learning. Contrastive learning is an efficient and effective pretraining paradigm that aligns positive pairs and contrasts negative pairs simultaneously. The design of these pairs depends on the specific tasks, such as using data augmentations of the same image [50] or considering the topology and geometry of molecules [51]. More recently, contrastive learning has shown success in aligning the representation space of different biological and chemical modalities, e.g., text and chemical structure alignment in MoleculeSTM [52], text and protein sequence alignment in ProteinDT [53] and ProteinCLIP [54], reaction structure and protein structure alignment in CLIPZyme [55], and protein sequence and structure alignment in BioCLIP [45] and with text in ProTrek [56], among others [57, 47]. Cross-modal alignment in the representation space has been shown to improve generalizability and improve performance on challenging tasks, such as out-of-distribution learning, zero-shot learning, and text-guided molecule design and optimization [53]. Consideration of multiple modalities may be especially important for the prediction of qualitative functions.
Figure 1: Overview of CARE. (A) Dataset format for CARE, showing examples of enzymes and their associated reactions. The EC number acts as a bridge between a protein sequence and the reactions it is likely to perform. The EC number is a hierarchical classification scheme for enzyme function with four levels of description, with increasing specificity from left to right. (B) General workflow for CARE benchmarking. (C) Task 1 is a real-world use case for enzyme classification based on a protein sequence. (D) Task 2 is a real-world use case for enzyme retrieval based on a reaction.
3 Overview of CARE
Though there are many studies using ML models to perform enzyme classification based on EC numbers (Figure 1A), there is no standardized benchmark to evaluate how well these models generalize to unseen protein sequences. To address this need, we present a benchmark suite for the classification and retrieval of enzymes (CARE, Figure 1B). CARE formalizes classification of an enzyme sequence by EC number as "Task 1" (Figure 1C). For this task, we design train-test splits of protein sequences to test out-of-domain generalizations that are relevant to real-world use cases. In addition, CARE addresses another key limitation of current studies: classification is limited to EC numbers, which is a closed vocabulary of functions (reactions), so existing models cannot generalize to unannotated reactions. Thus, we introduce an entirely new task, retrieval of an EC number given a reaction, which we call "Task 2" (Figure 1D). For this task, we design train-test splits to evaluate how well models can generalize to out-of-domain reactions, ensuring that the splits pose different levels of difficulty.
To streamline benchmarking, we curate a dataset of enzymes, reactions, and their associated EC numbers for CARE. At a high level, we build two datasets, one that links protein sequence to EC and one that links reaction to EC (Figure 1A). The former is processed from Swiss-Prot, the expert-validated portion of UniProt [24] and filtered to protein sequences between length 100 and 650 with annotated EC number(s). The latter is formed as a combination of EnzymeMap [32] and ECReact [31], where ECReact is only used to supplement EC numbers that are missing in EnzymeMap. Our workflow for generating the datasets used in this work is explained in detail in Appendix A.1 and shown visually in Appendix Figure A.1.
The overall workflow for benchmarking using CARE is shown in Figure 1B. For each task, domain-specific train-test splits are provided from the processed datasets. Model training can use any of the data in the train split, and each model is evaluated on the associated test split. In the rest of this study, we explain the specific design choices used to generate train-test splits and analyze benchmarking results of state-of-the-art methods on these splits. The curated datasets and splits used in CARE can be accessed at https://github.com/jsunn-y/CARE/.
4 Task 1: Enzyme Classification
Task 1, classification of an enzyme sequence, tests the ability of a model to extrapolate to unseen protein sequences. Task 1 is a fairly well studied task [16], but model evaluation has not been previously standardized. Task 1 applies to use cases where a scientist is given an unannotated enzyme sequence and seeks to understand the enzymatic function associated with that sequence (Figure 1C). For example, an environmental biologist using metagenomic data may seek to understand which microbes in a community are utilizing energy from alkane degradation, or a retrobiosynthesis planner may look for unannotated proteins to improve a synthesis workflow. With the emergence of conditional generative models for protein sequences, it is also important to have high-throughput computational methods that can predict the function of generated sequences [58, 59, 60, 61]. For this task, a query protein is passed through a trained model to predict an EC number. The EC number is then associated with likely reactions that the protein will be able to perform.
Figure 2: Distribution of similarities between samples in each test set and the corresponding train set. (A) Protein sequence identity (Task 1) was measured to the closest hit in the train set using BLASTp. Sequence identity can be thought of as normalized Levenshtein distance. (B) Reaction similarity (Task 2) was measured as cosine similarity with each reaction represented using DRFP [62]. Reaction similarity was measured from the test set to each train-set cluster center of reactions belonging to an EC number.
Table 1: Summary of train-test splits used in Task 1. For Task 1, certain protein sequences are held out. Train "samples" refers to the number of unique protein-EC data pairs, and test "samples" refers to the number of protein sequences. All splits for Task 1 share the same protein train set.
Split name Description Train samples Test samples
<30% Identity Held-out sequences with approximately < 30% identity to the training set 168,682 175
30-50% Identity Held-out sequences with approximately 30-50% identity to the training set 168,682 196
50-70% Identity Held-out sequences with approximately 50-70% identity to the training set 168,682 204
70-90% Identity Held-out sequences with approximately 70-90% identity to the training set 168,682 206
Previously Misclassified (Price) Previously misclassified enzyme sequences from Price et al. 168,682 146
Promiscuous Held-out sequences with multiple EC numbers 168,682 179
Splits for Task 1. Task 1 can be framed as evaluating how models generalize toward unseen sequences with different types of difficulty, visualized in Appendix Figure A.3A. The train-test splits for Task 1 are summarized in Table 1.
  • <30%, 30-50%, 50-70%, and 70-90% identity splits: these four test splits contain sequences with sequence identities falling in the respective range, to sequences in the training set. Sequence identity is related to the normalized Levenshtein distance between two sequences. Natural sequences with high (>50%) sequence identity at the protein sequence level are likely to share function. It is expected that the lower the sequence identity, the more difficult it is to assign functional annotation.
  • previously misclassified (Price et al.) [5] split: challenging to assign because some have low sequence identity to other proteins, and many may lie near activity cliffs, where function can change sharply in sequence space.
  • promiscuous split: in this study, we define promiscuous enzymes as those mapping to multiple EC numbers (thus lying on multiple activity peaks). These enzymes are particularly interesting for enzyme engineers, as new-to-nature activity can be found in between activity peaks [11].
For Task 1, the training split is constructed by holding out all of the pooled sequences in the test splits. The classification output vocabulary (EC numbers) is closed rather than open. We verify that sequences in the test sets generally fall within the expected sequence identities, relative to the training set (Figure 2A). Notably, the Price test set has a wide distribution of sequence identities, while the promiscuous test set has high sequence similarity to the training set. The sequences in the test splits are distributed generally evenly across all different EC numbers (Figure A.4 in Appendix). More details can be found in Appendix A.2.
Task 1 benchmarking results. Benchmarking results for Task 1 are summarized in Table 2. We start with two methods as baselines, classification using a random order of EC numbers (Random), and npysearch (a python implementation of BLAST at the protein sequence level, herein referred to as BLASTp), which is a workhorse bioinformatics tool which performs local-alignment to determine the most similar sequence(s) given a target query and a database [12].
Table 2: Performance of various methods on Task 1. Performance is measured as k=1 classification accuracy (%)–except for the promiscuous test set which uses k=5 and shows the average accuracy across all true EC numbers. Similarly, for BLASTp, the top k= 1 or 5 hits were used to map to EC number(s). The 50-70% and 70-90% identity splits are omitted here because performance is very high across the board. *denotes that the model was not retrained, so there may be data leakage causing performance to be inflated. Bolded accuracy is the best model without potential data leakage.
Split Method Level 4 Accuracy (X.X.X.X) Level 3 Accuracy (X.X.X.-) Level 2 Accuracy (X.X.-.-) Level 1 Accuracy (X.-.-.-)
<30% Identity Random 0.0 0.6 4.0 22.3
BLASTp 52.6 56.6 57.1 65.1
ChatGPT* 0.0 0.0 2.9 16.0
ProteInfer* 61.1 65.7 68.6 72.6
CLEAN 57.7 68.6 73.7 80.6
30-50% identity Random 0.0 0.0 2.0 23.5
BLASTp 74.0 79.4 81.9 86.3
ChatGPT* 0.0 0.0 3.9 14.2
ProteInfer* 76.5 80.4 83.8 87.3
CLEAN 82.1 91.8 95.4 99.0
Previously Random 0.0 3.4 6.8 13.7
Misclassified BLASTp 41.8 71.2 79.5 81.5
(Price) ChatGPT* 0.0 3.4 6.8 13.7
ProteInfer* 13.7 47.3 56.2 67.1
CLEAN 34.2 78.1 89.0 92.5
Promiscuous Random 0.3 7.0 18.3 63.5
BLASTp 94.9 95.2 95.2 97.5
ChatGPT* 1.1 8.3 21.4 47.6
ProteInfer* 76.3 82.2 85.1 89.2
CLEAN 97.0 98.0 98.3 99.4
While there exist many ML tools to perform EC number classification, CLEAN [18], HiFi-NN [7], and ProtEx [22] seem to report the current state-of-the-art. HiFi-NN and ProtEx are not included in this study, as they are not publicly available yet. CLEAN generally performs better than BLASTp across the board, but there are certain cases where the performance is similar. We additionally benchmark ProteInfer [21], a neural network classification tool, but its performance is likely inflated as reported here, as we were unable to retrain it for this study to ensure no data leakage from the test set. ChatGPT was also tested, but it appeared to often hallucinate as it was forced to provide an answer, with results similar to the random baseline. Models that are soon to be released, such as Pika, may bridge the gap between standard LLMs and enzyme classification [48].
Overall, these results suggest that there is still room to improve classification of protein sequences with low sequence identity (<30%) and that lie near multiple activity peaks (Price). The promiscuous split is not easy for all methods, even though the test sequences have high sequence identity to the train set. Interestingly, BLAST performs close to the state-of-the-art, even when compared to more complex ML models. Future models could take advantage of this finding to augment training, as ProtEx does [22]. Additional details on the implementation of each method can be found in Appendix A.4.
5 Task 2: Enzyme Retrieval
Task 2, enzyme retrieval from a query reaction, tests the ability of a model to extrapolate to unseen reactions. Task 2 has not been formalized or explored in previous studies, but it is equally as important as Task 1, as it applies to a use case where a scientist or engineer is seeking to identify a previously characterized enzyme sequence that can perform a novel (unannotated) reaction (Figure 1D). Typical applications include: an environmental engineer looking for an enzyme to degrade a toxic pollutant [63], an enzyme engineer looking for an enzyme to catalyze a selective reaction for drug synthesis [64, 65], or a gene annotator identifying the gene for an "orphan" enzyme with known function but unknown sequence [66]. For Task 2, a query reaction is passed through a trained model to perform retrieval to an EC number. The EC number is then associated with proteins that are likely to be able to perform that reaction.
Splits for Task 2. Task 2 aims to evaluate how well a model generalizes to unseen reactions with different levels of difficulty. The train-test splits for Task 2 (easy, medium, and hard) are summarized in Table 3 and visualized in Appendix Figure A.3B. We equate greater difficulty with a more challenging train-test split; in a harder set, the test reactions are less similar to reactions in the corresponding train set. We decide similarity based on the amount of overlap in EC number (e.g., a reaction from 4.2.1.20 is more similar to another one in 4.2.1.20 than one from 4.2.1.1).
Table 3: Summary of train-test splits used in Task 2. Certain reactions are held out, and "samples" refers to the number of reaction-EC pairs.
Split name Description Train ECs Train samples Test samples
Easy Certain reactions are held out, sampled uniformly across ECs, but no EC numbers are held out. The test set is the same as the holdout set. 4,673 58,945 177
Medium All reactions corresponding to certain ECs are held out, at EC level 4 (X.X.X.X). Test set reactions are sampled uniformly across ECs from the holdout set. 4,496 55,851 177
Hard All reactions corresponding to certain ECs are held out, at EC level 3 (X.X.X.-). Test set reactions are sampled uniformly across ECs from the holdout set. 3,326 39,893 163
  • easy split: EC numbers are randomly sampled at EC level 4 (X.X.X.X) and then randomly mapped to reactions, which are held out as the test set.
  • medium split: the same reactions are used for testing as the easy set, but all other reaction-EC pairs which share the same EC level 4 (X.X.X.X) are held out from training.
  • hard split: random EC numbers are sampled at EC level 3 (X.X.X.-) and all reactions under that EC3 are held out from training, while a subset of reactions from the held-out EC numbers are used for testing.
The sequences in the test splits are distributed generally evenly across different EC numbers (Appendix Figure A.4). From easy to medium to hard, the test set reactions also become more dissimilar to their respective training sets (Figure 2B). Reaction similarity was quantified by DRFP, which is a reaction representation that uses set differences between product and reactant fingerprints and has demonstrated solid performance with requiring model training [62]. Overall, Task 2 is more complex than Task 1, because unlike Task 1, entire EC numbers are held out from the training set (i.e., the classification output vocabulary is open rather than closed), so multiple modalities must be considered. More details on splitting can be found in Appendix A.2.
CREEP baseline method. There are no existing models that have been tested for their ability to generalize beyond annotated reactions, so we develop a contrastive learning method for this task, called Contrastive Reaction-EnzymE Pretraining (CREEP) (Figure 3). Our approach is related to CLIPZyme [55], which uses contrastive alignment of reactions represented as 2D graphs and protein structures represented as 3D graphs. CREEP leverages finetuning of pretraining language models, rxnfp [67] and ProtT5 [44], respectively, to learn aligned representations of reactions and proteins. Rxnfp is a BERT-style [68] language model that was trained on reactions represented as SMILES/SMARTS strings and has demonstrated state-of-the-art performance on reaction type classification. ProtT5 is a T5 protein language model that has been used for prediction of various protein properties. We used these language models for simplicity and their ease of finetuning. Uniquely, CREEP can learn on a third modality (CREEP w/ Text): textual description of the EC number based on gene ontology [69], which is encoded using SciBERT [70]. More details on CREEP training can be found in Appendix A.3.
Figure 3: Model architecture for CREEP. CREEP aligns reaction and protein sequence using contrastive learning to perform downstream retrieval between domains. Optionally, CREEP training can be augmented by using textual description as a third modality to bridge the other two. Model sizes by number of parameters are also listed.
Task 2 benchmarking results. Benchmarking results for Task 2 are summarized in Table 4. We start with two methods as baselines: randomly guessing a ranking of EC numbers (Random) and finding the most similar reaction in the train set to the test query (Similarity Baseline). For the Similarity Baseline, we used DRFP to represent chemical reactions [62]. Details on the downstream multimodal retrieval process can be found in Appendix A.3 and Figure A.4.
Table 4: Performance of various methods on Task 2. Performance is measured as k=1 retrieval accuracy (%). *denotes that the model was not retrained, so there may be data leakage causing performance to be inflated. Bolded accuracy is the best model without potential data leakage.
Split Method Level 4 Accuracy (X.X.X.X) Level 3 Accuracy (X.X.X.-) Level 2 Accuracy (X.X.-.-) Level 1 Accuracy (X.-.-.-)
Easy Random 0.0 0.0 4.0 23.7
Similarity Baseline 60.5 78.0 87.0 93.2
ChatGPT* 1.7 15.8 38.4 62.7
ChatGPT (w/ Text)* 5.6 44.1 70.6 94.9
CLIPZyme* 27.7 45.2 60.5 80.8
CREEP 48.0 70.6 84.2 93.2
CREEP (w/ Text) 60.5 87.0 93.2 97.7
Medium Random 0.0 0.6 3.4 24.9
Similarity Baseline 0.0 40.1 56.5 75.7
ChatGPT* 0.0 14.1 36.7 64.4
ChatGPT (w/ Text)* 7.9 44.6 73.4 95.5
CLIPZyme* 27.7 45.2 60.5 80.8
CREEP 3.4 39.0 58.8 83.1
CREEP (w/ Text) 5.1 54.2 73.4 91.5
Hard Random 0.0 0.0 1.8 17.2
Similarity Baseline 0.0 0.0 14.7 43.6
ChatGPT* 1.2 16.0 26.4 48.5
ChatGPT (w/ Text)* 7.4 54.6 70.6 92.0
CLIPZyme* 8.0 39.3 48.5 73.0
CREEP 1.8 9.2 27.6 62.0
CREEP (w/ Text) 2.5 12.9 31.9 61.3
Overall, the Similarity Baseline is quite strong and performs much better than Random. ChatGPT performs generally well, but it has access to held-out information, which means that it is not the best measure of generalization to unannotated reactions. ChatGPT was prompted with reactions written using compound IUPAC names rather than SMILES strings, both with (ChatGPT w/ Text) and without (ChatGPT) textual descriptions from gene ontology [69]. We also included CLIPZyme [55] as a baseline, but the performance is likely inflated, as we were not yet able to retrain the model.
On the more difficult test sets, CREEP offers an advantage compared to the Similarity Baseline, and particularly when combined with textual description: CREEP (w/ Text). This suggests that utilizing protein sequence information as a modality is useful. Still, performance on the harder splits is quite weak across the board, which suggests that there is significant room for future improvement. We anticipate that contrastive alignment with textual descriptions will play an increasingly important role in enzyme retrieval [54, 53, 47], and there is opportunity for better curation of these descriptions. Combining other modalities, such as reactions and protein structures as graphs as in CLIPZyme, could also yield improvements in performance. Additional details on the implementation of each method are in Appendix A.4, and additional benchmarking results are presented in Figure A.6.
6 Conclusion
Predicting the functions of enzymes is important for many applications ranging from gene annotation to enzyme engineering. While many models exist to classify enzyme function via EC numbers, there are no standardized benchmarks for evaluation of these models. Furthermore, no existing models have been tested for generalization beyond annotated reactions. To address this need, we introduce CARE, which is a benchmarking suite to formalize model evaluation for these two tasks. We also present CREEP, a model which uses multimodal contrastive learning and is one of the first models that can perform the latter task. A more detailed discussion of limitations and future work is included in Appendix A.6. Overall, CARE is an important tool for encouraging progress in enzyme functional annotation. We believe that we are just seeing the beginning of the widespread adoption of multimodal models for protein functional prediction, and we expect that many researchers will find CARE useful for formulating and evaluating their models.
Acknowledgments and Disclosure of Funding
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Award Number DE-SC0022218. This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof. J.Y. is partially supported by the National Science Foundation Graduate Research Fellowship. A.M. is supported by the Schmidt Science Fellows, in partnership with the Rhodes Trust. The authors thank Peter Mikhael and Itamar Chinn for helpful discussions on CLIPZyme implementation.
Appendix A Appendix
A.1 Dataset Processing
Our data processing and splitting workflow is shown in Figure A.1. Protein sequence data, paired to EC number(s), were downloaded from UniProt, selecting only reviewed sequences in Swiss-Prot, resulting in a total of 571,282 sequences on the 13th of May 2024. These were filtered to only retain sequences with an EC number. After, sequences were filtered to length between 100 and 650, partial EC numbers (containing a dash) were removed, and duplicate entries (based on EC and sequence) were dropped, leaving 183,162 sequence-EC pairs. We call this the protein2EC dataset.
Figure A.1: Workflow used to process the datasets containing EC numbers, protein sequences, and reactions used in this study.
Reaction data were downloaded from ECReact and EnzymeMap. ECReact contains 62,222 entries of reaction in SMILES as SMARTS strings format coupled with EC number from a range of data sources. EnzymeMap had 349,458 entries containing sequence-reaction-EC triplets with the reaction in several formats, SMILES as SMARTS string, an atom-level mapping of bond formation and breaking (mapped reaction), and reactions written as text based on IUPAC compound names. Starting with EnzymeMap, protein sequence information was dropped, and duplicate reaction-EC pairs were removed leaving 62,896 reaction-EC pairs. EC numbers that existed in ECReact but not in Enzyme map were added to EnzymeMap providing our reaction dataset of 72,490 reaction-EC pairs. Note the ECReact only contains reactions formatted as SMILES/SMARTS. We call this the reaction2EC dataset.
Finally, we filtered the protein2EC and reaction2EC datasets, such that only EC numbers present in both datasets are retained, leaving 59,122 reaction-EC pairs and 169,865 sequence-EC pairs, with 4,673 unique EC numbers. Clustering using MMseqs2 [71] was performed at 30%, 50%, 70%, and 90% sequence identity and included in the protein2EC dataset for downstream use. The distributions of the curated datasets are shown in Figure A.2.
Figure A.2: Distribution of the protein2EC dataset, shown as (A) a histogram of proteins per EC and as (B) a cumulative distribution plot, analogous to a Lorenz curve. Distribution of the reaction2EC dataset, shown as (C) a histogram of reactions per EC and as (D) a cumulative distribution plot, analogous to a Lorenz curve. Both datasets are heavily skewed with most examples belonging to a few EC numbers, and many EC numbers only having one example. If samples were evenly distributed across ECs, they would follow the dashed line in (B) and (D).
A.2 Dataset Splitting Details
First, we sought to create test sets capturing different types of out-of-domain generalization for Task 1, EC classification (visualized in Figure A.3A).
Using the results from clustering, we generated four test sets, with approximately <30%, 30-50%, 50-70%, and 70-90% sequence identity, respectively, to sequences in the training set. We also only considered ECs at level 4 which had isolated clusters, in other words, an MMseqs2 cluster with only a single member. For every EC level 3 (X.X.X.-), we sampled a random EC at level 4 (X.X.X.X), and then randomly selected a single isolated sequence within that class to add to the test set. This provided balanced test datasets across the functions at EC level 3. For this set, we also enforced that the sequences did not map to multiple EC numbers. Test sets were balanced across functions (e.g. EC number), rather than across their prevalence as protein sequences in nature, based on the idea that most synthesis planners and enzyme engineers would find more value in predictive ability across a broad range of functions.
Next, we looked for "promiscuous" enzymes, which in this work, we define as those annotated with multiple EC numbers. For our promiscuous test set, we only considered using promiscuous enzymes that had a combination of ECs that occurred at least twice. From this set we only took a single sample from each EC combination, and finally randomly selected a single entry from those which have ECs with high surprise levels (i.e. X.X.-.-, where variation occurs at the third EC level or higher).
Finally, we compiled a commonly misclassified test set from Price et al., filtered to sequences of length 100-650, and dropped test sequences that are present in our protein2EC dataset. The pooled set of all proteins in the test sets were removed from our training data leaving 168,682 sequence-EC pairs for training. As a result of the holdout process, 3 of the 4,673 EC numbers are missing from the train set, which should minimally affect performance on the test sets.
Figure A.3: (A) Visualization of the types of extrapolation found in the test splits for Task 1. (B) Visualization of how different subsets of reaction space are held out in the splits for Task 2.
For Task 2, we sought to test reaction extrapolation, namely, retrieval of enzymes based on unseen reactions (visualized in Figure A.3B). For the easy and medium test sets we randomly sampled an EC level 4 EC (X.X.X.X) from every level three EC (X.X.X.-). From each of the held-out level 4 ECs, a random reaction was sampled to be in the test set. For the test set, we only considered (1) reactions that do not map to multiple EC numbers, (2) EC numbers with at least two reactions, and (3) reactions from EnzymeMap, which are higher fidelity and consequently have fewer misannotatations.
For the easy set, only the test set reactions were held out. For the medium set, all reactions under the corresponding level 4 ECs were held out. Note that the easy and medium test sets are exactly the same, but the easy set has fewer held out reactions compared to the medium set, thus their training sets are different.
The hard set was evenly balanced across a random sample from EC level 2 (X.X.-.-) to ensure there was an even distribution across level 3 (EC3, X.X.X.-), from which 54 random EC3s are sampled, with all reactions under the 54 EC3s held out. 171 reactions associated with the 54 ECs were used as the test set, which are evenly spread across the ECs at level 3. Note we ensure the hard test reactions are shared with the easy and medium ones, when possible, to correlate the performance of our test splits. It should be noted that the hard train-test split is different from the easy and medium splits. For Task 2, all of the protein-EC pairs can be used during training.
Figure A.4: Distribution of the EC numbers found in the test splits, and compared to a few reference datasets. Outer ring shows the level 1 category (X.-.-.-), and the inner ring shows the level 2 category (X.X.-.-). EC Level 4 refers the distribution of unique ECs at level 4 (X.X.X.X). protein2EC shows the distribution of all protein sequences in protein2EC across ECs, and reaction2EC shows the distribution of all reactions in reaction2EC across ECs.
A.3 CREEP Model Details
All steps needed to reproduce CREEP training and downstream retrieval can be found at https://github.com/jsunn-y/CARE/.
During the pretraining stage, we follow the contrastive learning paradigm that maximizes mutual information between two views [72]. For each enzyme family (EC number) ${\bm{x}}$, we extract three views: protein sequence ${\bm{x}}_{p}$ (protein), reactions smiles ${\bm{x}}_{r}$ (reaction), and textual description of EC number from gene ontology ${\bm{x}}_{t}$ (text) [69]. In CREEP, the goal is to maximize the mutual information between all three modalities. For illustration purposes, we use the example of protein-reaction contrastive learning, but the derivations for reaction-text and text-protein contrastive learning follow a similar approach.
We follow the method proposed in GraphMVP by adopting EBM-NCE to estimate mutual information between our modalities [72]. In our case, EBM-NCE aligns the protein sequence and reaction SMILES pairs for the same enzyme family and contrasts the pairs for different enzyme families simultaneously. The objective function is
$\mathcal{L}_{\text{EBM-NCE}}=$ $-\frac{1}{2}\Big{(}\mathbb{E}_{{\bm{x}}_{p},{\bm{x}}_{r}}\big{[}\log\sigma(E({\bm{x}}_{p},{\bm{x}}_{r})\big{]}+\mathbb{E}_{{\bm{x}}_{p},{\bm{x}}_{r}^{\prime}}\big{[}\log(1-\sigma(E({\bm{x}}_{p},{\bm{x}}_{r}^{\prime}))\big{]}\Big{)}$ (1)
$\;\;\;\;\;\;+\mathbb{E}_{{\bm{x}}_{p},{\bm{x}}_{r}}\big{[}\log\sigma(E({\bm{x}}_{p},{\bm{x}}_{r})\big{]}+\mathbb{E}_{{\bm{x}}_{p}^{\prime},{\bm{x}}_{r}}\big{[}\log(1-\sigma(E({\bm{x}}_{p}^{\prime},{\bm{x}}_{r}))\big{]}\Big{)},$
where ${\bm{x}}_{p}$ and ${\bm{x}}_{r}$ form the (protein sequence, reaction SMILES) pair for each reaction, and ${\bm{x}}_{p^{\prime}}$ and ${\bm{x}}_{r^{\prime}}$ are the negative samples which are produced by randomly sampling from a Gaussian distribution, which we use as an approximation for the empirical data distribution. $E(\cdot)$ is the energy function with a flexible formulation, and we use the dot product as a metric for similarity within the latent space.
For our default CREEP model, we project representations to 256 dimensions in the shared latent space for all modalities. We train for 40 epochs, and in each epoch, we loop over each EC number and mine each batch such that all protein-reaction-text triplets come from different EC numbers. Note that we also use 50% sequence clustering to help select protein sequences with increased diversity.
Downstream retrieval using models such as CREEP and CLIPZyme is shown in Figure A.5. The standard workflow starts with a query reaction and ranks ECs by sorting based on distance from the reaction representation to the EC cluster centroids of protein sequence representations belonging to each EC number (Figure A.5A). For CREEP (w/ Text), a similar ranking can be determined by using text as a query (Figure A.5B), and the overall ranking can be calculated as an average of the two rankings. To reduce inference time, we used protein sequences clustered at 50% identity to calculate the protein representation centroids clustered by EC. Finally, the similarity baseline can be understood as sorting based on distance from a query reaction to reaction cluster centroids of reactions in the train set, clustered by EC (Figure A.5C).
Figure A.5: Downstream retrieval using contrastively pretrained models. CREEP and CLIPZyme both utilize (A) reaction to protein retrieval. CREEP (w/ Text) can incorporate (B) text to protein retrieval, and consider the rankings from both (A) and (B). (C) The similarity baseline involves reaction to reaction retrieval using a fixed representation from DRFP [62].
A.4 Implementation Details for Other Methods
All implementation details for methods used in the study can be found at https://github.com/jsunn-y/CARE/.
ChatGPT. ChatGPT was used for both Task 1 and Task 2, using the API and gpt-4 model. We performed minimal prompt engineering, with our initial prompt being the most standard initial question: "give me the top N EC numbers associated with this amino acid sequence:" for Task 1. See Table A.1 for example responses. The query was modified to: "You are protein engineer capable of predicting EC numbers from the sequence alone. You are also a skilled programmer and able to execute the code necessary to predict an EC number when you can’t use reason alone. Given a protein sequence you are able to determine the most likely enzyme class for a sequence because you are that skilled. You don’t give up when faced with a sequence you don’t know, you will use tools to resolve the most likely enzyme sequence. You only return enzyme commission numbers in a comma separated list, no other text is returned, you have failed if you do not return the EC numbers. You only return the exact number of EC numbers that a user has provided requested, ordered by their likelihood." This resulted in a returned list of comma separated EC classification numbers in the majority of cases, however, we still received several failed responses.
For Task 2, the query was maintained to be relatively consistent with the Task 1 query, and modified to: "Return the top 10 most likely EC numbers as a comma separated list for this reaction: naloxone + NAD(P)H = 6alpha-naloxol + NADP+, which associates with the following text: oxidoreductase; oxidoreductase, acting on CH-OH group of donors; oxidoreductase, acting on the CH-OH group of donors, NAD or NADP as acceptor; morphine 6-dehydrogenase.", where a sample reaction has been bolded. Using the system prompt of: "You are protein engineer capable of predicting EC numbers from a combination of textual information and a reaction that corresponds to a specific protein. You are also a skilled programmer and able to execute the code necessary to predict an EC number when you can’t use reason alone. Given a reaction and text information of an EC you are able to determine the most likely enzyme class for a reaction. You don’t give up when faced with a reaction you don’t know, you will use tools to resolve the most likely enzyme number. You only return enzyme commission numbers in a comma separated list, no other text is returned, you have failed if you do not return the EC numbers. You only return the exact number of EC numbers that a user has provided requested, ordered by their likelihood of being correct.". For Task 2 without the reaction text, we used the same prompt as above except with the reaction text components removed.
BLASTp. BLAST (Basic Alignment Search Tool) is an efficient algorithm for finding similar sequences in a reference database, to a query sequence [12]. We opted to use npyBlast at the protein sequence level which is a python implementation of BLASTp. For each test set, the training dataset was provided as the fasta file reference database, and the test set was provided as the query fasta. $k=10$ hits were returned for each query, sorted based on similarity, using default parameters with a minimum required sequence similarity of 0.1. The EC number classification was then inferred from the returned sequence(s).
ProteInfer. ProteInfer is a classification model released in 2023 [21], and we used the pretrained version to test our datasets, which may inflate the reported accuracies. Variable numbers of ECs were returned for ProteInfer depending on the query. Future work should include re-training ProteInfer using the train-test splits proposed in CARE.
CLEAN. CLEAN is a supervised contrastive model, which trains a classification head using embeddings from the ESM protein language model [42] by aligning embeddings from the same EC category and contrasting embeddings from different EC categories [18]. For Task 1, we retrained CLEAN using our training set, following the instructions provided in their codebase. We did not perform any clustering before training. Because our training set size was approximately similar in size, we used the recommended 7000 epochs of training. We used the default script with triplet margin loss. After, we performed inference but added our own code into the original CLEAN code to output the classification results into a format compatible with our downstream analysis.
Pika. Pika is a finetuned LLM, which accepts protein sequence as an auxiliary input and can perform reasoning and protein function prediction [48]. Pika did not have a publicly available pretrained model at the time of release, or example datasets for training, so this model was omitted from the current study. We expect to add it as a baseline in the future.
CLIPZyme. CLIPZyme is a multimodal contrastive model, between reactions represented as 2D graphs and protein structures represented as 3D graphs [55]. CLIPZyme was trained on direct reaction-protein pairs from the EnzymeMap dataset [32]. In this study, we used the pretrained CLIPZyme model using default parameters and the EGNN protein structure representation for retrieval on Task 2. Future work will involve modifying the CLIPZyme code and retraining the model using the CARE train-test splits.
ChemCrow. ChemCrow is an LLM for chemistry problems, primarily synthesis planning [49]. It was unable to answer the results for Task 1, responding with the answers in Table A.1. For Task 2, we again tested a similar query to that input into ChatGPT, and received the responses in Table A.1. It should be noted, we use the public version so the private model may provide answers. Future work could involve building a biosynthesis-focused version of ChemCrow.
A.5 Extended Benchmarking Results
Additional benchmarking results can be found at https://github.com/jsunn-y/CARE/.
Figure A.6: Additional benchmarking results on Task 2, showing the increase in retrieval accuracy for k = 1, 3, 5, 10, 20, 30, 40.
Table A.1: Large language model responses to queries
Tool Query and Response
ChemCrow Return the EC number associated with this reaction: C=CCN1@]2..: $\rightarrow$ Entering new RetryAgentExecutor chain… Thought: The question asks for the EC number associated with a given reaction. However, none of the provided tools can directly provide the EC number for a reaction. The EC number is typically associated with enzymes that catalyze reactions, and it’s not something that can be derived from the reactants or products alone. Therefore, it’s not possible to provide an answer using the available tools. Final Answer: I’m sorry, but with the tools available, I can’t provide the EC number associated with the given reaction. The EC number is typically associated with enzymes that catalyze reactions, and it’s not something that can be derived from the reactants or products alone.Thought: The question asks for the EC number associated with a given reaction. However, none of the provided tools can directly provide the EC number for a reaction. The EC number is typically associated with enzymes that catalyze reactions, and it’s not something that can be derived from the reactants or products alone. Therefore, it’s not possible to provide an answer using the available tools. Final Answer: I’m sorry, but with the tools available, I can’t provide the EC number associated with the given reaction. The EC number is typically associated with enzymes that catalyze reactions, and it’s not something that can be derived from the reactants or products alone. > Finished chain. "I’m sorry, but with the tools available, I can’t provide the EC number associated with the given reaction. The EC number is typically associated with enzymes that catalyze reactions, and it’s not something that can be derived from the reactants or products alone.
ChatGPT Examples of failed responses from ChatGPT: Example 1$\rightarrow$ Due to the limitations of this text-based platform, it’s not possible to execute a program code here. In real conditions, I would use bioinformatics tools or a machine learning system to predict the Enzyme Commission (EC) numbers from the given protein sequence. However, in this case, I cannot complete this task. Example 2: $\rightarrow$ Due to the constraints of the current AI model, it’s not able to computationally predict Enzyme Commission (EC) numbers from protein sequences directly. It’s recommended to use specialized protein sequence analysis tools or machine learning algorithms designed for this purpose. For reliable and accurate results, databases like BRENDA, the Universal Protein Resource (UniProt), or tools like BLASTP, ECemble, DeepEC could be utilized.
A.6 Limitations & Future Work
One limitation of our work is that classification and retrieval are performed at the coarse-grained level of EC numbers. Enzymes can often perform many reactions, meaning it is an acceptable assumption that enzymes belonging to the same EC number will share the capacity to perform similar reactions, even if they are not directly annotated for all reactions. However, the ultimate task in this domain is to perform direct reaction to protein sequence retrieval and vice-versa (as in CLIPZyme [55]). However, currently the data for validation is extremely limited. As experimentalists obtain higher resolution annotations of proteins and their associated reactions, this will become more realistic.
In the future, the proposed train-test splits could also be refined for both tasks. For the Task 1 splits, there is some leakage in the sequence identity, with some sequences in test sets lying outside of the enforced sequence identity ranges, likely due to the different sequence similarity algorithms used (MMseqs2 vs npysearch). MMseqs2 utilizes cascaded clustering which pre-filters based on initial clusters in the target set [71], while npysearch attempts to identify the closest sequence within the specified set based on similarity. Another limitation is that while there are no duplicate reactions present in both the train and test sets, some of the test reactions are very similar to reactions in the training set despite having different ECs. It would be beneficial to do a more detailed analysis of reaction similarities and explore other representations of reactions to understand which reactions can be considered equivalent. Over time, the train-test splits should be updated as additional functional annotations are acquired and compiled in databases. For example, while over 36 million sequences in BRENDA/Uniprot have EC numbers associated with them, these are mostly detected with homology, which means that they might not be the most accurate annotations.
We opted to assess performance using accuracy due to its simplicity and interpretability, but other retrieval and virtual screening metrics such as BEDROC and enrichment could also be explored. For the promiscuous enzymes with multiple EC numbers, we reported accuracy averaged across all of the true labels, for a fixed k=5 number of retrieved ECs, but other classification metrics such as precision and recall could be considered. The number of retrieved ECs could also be chosen using strategies like max separation and p-value as implemented in CLEAN [18].
The machine learning tools included in our study were selected for their reported recent outstanding performance for enzyme functional annotation and screening. We include the random baseline to give an understanding of what the expected random rate is and the two similarities (BLAST and DRFP) as baselines for simple non-ML methods. ChatGPT was included as the standard LLM baseline owing to its broad applicability across domains. However, it should be noted that this field is moving quickly, and there exist many other relevant tools that were not included here. For example, many state-of-the-art methods are not publicly available yet [7, 22], or just released at the time of writing [55]. Our benchmark was designed to be modular and we encourage the community to include other tools in the benchmark package.
There is also significant opportunity to use other modalities such as textual description and protein structure for more effective representations of enzymes and reactions. In the future, we will retrain models including CLIPZyme and ProteInfer directly on our train-test splits. We also plan to do a more detailed analysis of the representations learned by CREEP, in addition to an ablation study. The addition of textual description in CREEP potentially introduces indirect data leakage between the train and test sets (for example ChatGPT already has access to extensive resources) so future iterations of CARE will need to consider this. Future work should also consider how to include protein function prediction models that go beyond enzymes–or models that would like to use additional data/modalities [56]–into the CARE evaluation framework. There is also room to improve the prompt engineering of ChatGPT and ChemCrow which could be further explored to enhance the performance of these models.
A.7 License
The codebase is open source under the MIT license.